# A Comparative Study of Image Restoration Networks for General Backbone Network Design

### 24 Jan 2024 v2

@@ Abstract
Abstract
In this supplementary material, we present additional experiments and results to complement the main manuscript. First, we illustrate the importance of a general backbone network for multi-task optimization on all-in-one restoration. Then, we complement details about schematic diagrams of mentioned three kinds of architectures. Next, we provide complete benchmark results for the five image restoration tasks and present visual results of X-Restormer with other benchmark networks. After that, we conduct comprehensive ablation study to further verify the effectiveness of our X-Restormer. Finally, we provide a model complexity comparison of our X-Restormer with Restormer.
@@ 1 Introduction
Image restoration aims to generate high-quality images from degraded images. In recent years, deep learning has achieved great success in this field, with numerous networks being proposed to address various image restoration tasks. Initially, networks are primarily designed to solve specific restoration tasks and are typically validated only on selected tasks. As deep learning techniques have continued to evolve, there has been an increasing focus on the development of general-purpose networks that can be applied to a broad range of tasks. This trend is particularly evident in the high-level vision field, where new backbone networks are being designed to support multiple tasks [17, 30], including classification, detection and segmentation. For image restoration, although more and more restoration backbones can handle multiple restoration tasks, their task generality is still limited, as illustrated in Figure 2. For instance, SwinIR [25] achieves state-of-the-art performance on image super-resolution (SR) but falls short on image deblurring and dehazing. Conversely, Restormer [50] performs exceptionally well on image dehazing and deraining but is less effective on image SR. This discrepancy can be attributed to the fact that the characteristics of image degradation vary across different image restoration tasks. While all image restoration tasks involve mapping degraded images to clean images, the requirements for the capability of backbone networks differ depending on specific tasks.
@@
Figure 1: Relative performance difference of different backbone networks on five image restoration tasks [^2] We set the minimum average performance of the networks on test sets in Table 2 for the task (i) as the lower bound $P^{(i)}_{lower}$, and the performance of X-Restormer as the upper bound $P^{(i)}_{upper}$. The ordinate of each point with performance $P^{(i)}$ is calculated as $(P^{(i)}-P^{(i)}_{lower})/P^{(i)}_{upper}$.. The existing representative networks exhibit diverse performance on these tasks, while our method presents superior task generality.
@@
Designing a general image restoration backbone network presents a significant challenge. However, the development of such a network holds considerable value, as it has the potential to greatly reduce costs associated with research and application. To achieve this goal, we first conduct a comparative study of mainstream backbone networks on the representative tasks, including image SR, denoising, deblurring, deraining and dehazing. These five tasks are chosen due to the distinct characteristics of their degradation. Five representative backbone networks are selected in the study, including MPRNet [49], Uformer [44], SwinIR [25], Restormer [50] and NAFNet [6]. These five networks encompass classic architectures such as U-shape architecture, plain residual-in-residual architecture and multi-stage progressive architecture. They also employ several common operators, including convolution, spatial self-attention and transposed self-attention [50]. We benchmark the five representative methods on the selected five tasks. The experimental results cleaårly reflect the performance disparity of different backbone networks on different tasks. We then conduct a detailed analysis of the characteristics of these tasks and these backbone networks to explain the reasons behind the performance differences. Based on the comparative study, we believe that a general backbone network must be highly comprehensive in terms of functionality that meets the diverse needs of various tasks.
@@
It is noteworthy that Restormer stands out in the comparative study, ranking within the top two across all five tasks. This superior performance can be attributed to several key designs. First, Restormer’s U-shape architecture allows it to process large-size inputs. Then, the network employs transposed self-attention, utilizing channel-wise feature maps as tokens, which enables mapping with a global receptive field. Additionally, the incorporation of numerous depth-wise convolutions activates the spatial information interaction. From a functional perspective, Restormer integrates the capabilities of the other compared networks, thereby exhibiting commendable task generality. However, the spatial mapping ability of Restormer appears to be somewhat deficient, as indicated by its quantitatively and qualitatively subpar performance in comparison to SwinIR for SR.
@@
This inferiority is hypothesized to originate from the inherent challenge of detail reconstruction posed by the U-shape architecture, coupled with the relatively weak spatial mapping capability of depth-wise convolution, particularly when compared to spatial self-attention (i.e., window-based self-attention in SwinIR). To address this limitation, a plausible solution is the introduction of spatial self-attention into Restormer. To validate the effectiveness of our design, we alternately replace half of the transposed self-attention block with the overlapping cross-attention block [8] to build a new network, X-Restormer. This simple modification significantly enhances the performance of Restormer without increasing the number of parameters. Experiments reveal that X-Restormer obtains state-of-the-art performance on all five tasks, exhibiting the best task generality.
@@
    We conduct a comparative study by constructing an image restoration benchmark, highlighting the challenges faced by existing backbone networks in task generality.
@@
    Based on the benchmark results, we perform a detailed analysis of the characteristics of different degradations and different networks. We emphasize that the general backbone network design must meet the functional requirements of diverse restoration tasks.
@@
    By further enhancing the spatial mapping ability of Restormer, we design a preliminary general backbone network, X-Restormer. Without additional parameters, X-Restormer achieves significant performance improvement and exhibits superior task generality.
@@ 2 Related Work
Image restoration networks.
Since SRCNN [12] first introduces deep learning to image SR and obtains superior performance over conventional methods, numerous deep networks have been proposed for various image restoration tasks such as image SR [26, 60, 8], denoising [54, 55, 44], deblurring [1, 5], deraining [48, 4, 29] and dehazing [46, 41, 40]. In earlier years, most deep networks are designed for specific image restoration tasks. For instance, [59] uses the channel attention mechanism to construct a CNN-based SR network. [56] builds a denoising network by inputting multiple downsampled sub-images and a noise level map. [15] utilizes residual connections for a deraining network. [3] estimates the transmission rate and inverts the atmospheric scattering model using a CNN to achieve image dehazing. [34] designs a multi-scale network for image deblurring.
Figure 2: Selected five image restoration tasks. The degradation in each task varies significantly.
Recently, more and more methods have been developed to tackle multiple image restoration tasks. [49] builds a multi-stage CNN for deraining, deblurring and denoising. [44] designs a U-shape Transformer for deraining, deblurring and denoising. [25] implements a Swin Transformer-based network that achieves state-of-the-art performance on SR, denoising and compression artifact reduction. [50] proposes a novel transposed self-attention to build a U-shape network for deraining, deblurring and denoising. [6] constructs a U-shaped CNN for denoising and deblurring.
Difference from the previous network design research.
While previous works have proposed networks that excel in various image restoration tasks, their primary focus is on constructing stronger networks to achieve performance breakthroughs on specific tasks. In contrast, this work pays more attention to the task generality of the backbone network, possessing a vision different from previous works. More specifically, our objective is to explore the design principles and directions of general image restoration networks. We are not seeking to create powerful networks for peak performance on a single or some specific tasks, but rather to ensure satisfactory performance across a diverse range of tasks. Regarding the concrete implementation, we do not intend to construct complex network architectures or modules. Our preference, rather, is to enhance task generality through the use of the simplest methodology available.
@@ 3.1 Overview of Image Restoration Tasks
We select five representative tasks for the benchmark experiments. These tasks, exemplified in Figure 2, are chosen based on two primary reasons. First, they are very common restoration tasks. Second, and more importantly, the degradation characteristics of these tasks vary significantly. As such, they can provide a robust means of evaluating the task generality of an image restoration backbone network.
@@
where $f_{motion}^{t}(\cdot)$ represents the motion function under different continuous exposure times. This degradation is related to the global motion offset of the image. Therefore, the ability to utilize long-range information and even global information is important for deblurring networks.
@@
where $R$ denotes the additive rain streak, simulated by the physics models, such as [23, 28]. The difference between this degradation and Gaussian noise is that the added $R$ is not evenly distributed on the image and has a correlation with the image content.
@@
where $t(\cdot)$ represents the transmission function and $t(I_{GT})$ is associated with the distance from the scene point to the camera. This degradation is intrinsically linked to the depth information within the image. Consequently, the incorporation of global information is important for the effective functioning of dehazing networks.
@@ 3.2 Characteristics of Typical Backbone Networks
The architectures of mainstream image restoration networks can be broadly classified into three categories: U-shape encoder-decoder, plain residual-in-residual and multi-stage progressive. Schematic diagrams of these architectures are provided in Supp. The U-shaped encoder-decoder architecture performs down-sampling and up-sampling operations on features, enabling networks to handle features of varying scales. This architecture allows networks to accept large-size input, and the effective receptive field of the network expands rapidly with down-sampling. Typical U-shape networks include Uformer [44], Restormer [50]. The multi-stage architecture divides the entire network into several sub-networks and progressively processes features, which are primarily used for image deraining and deblurring. Common networks based on this architecture include MPRNet [49] and HINet [5]. The plain residual-in-residual architecture is composed of several residual groups, each of which consists of several residual blocks. This architecture maintains the original size when processing features, which is favorable for the reconstruction of high-frequency information, but it comes at a high computational cost. Typical networks include RCAN [59] and SwinIR [25].
The core operators for constructing an image restoration network can be mainly categorized into three types: convolution, spatial self-attention and transposed self-attention. These operators are shown in Figure 3. The convolution calculates a fixed-size filter and processes the entire feature map through a sliding window, which is the major component of many networks, such as RDN [62]. Spatial self-attention is typically implemented as window self-attention in image restoration tasks. It calculates the attention matrix within a fixed window size, generating content-aware weights that are functionally similar to a large kernel dynamic filter. This operator has strong local fitting ability and shows superior advantages on SR and denoising [7]. Transposed self-attention treats the entire feature of each channel as a token to calculate the attention matrix on the channel dimension. This operator directly deals with global features, and when combined with depth-wise convolution, it shows remarkable performance in multiple restoration tasks [50].
The selected five representative backbone networks for the benchmark experiment encompass the abovementioned architectures and core operators, as presented in Table 1.
Table 1: Architectures and core operators of the five selected backbone networks for the image restoration benchmark.
@@ 3.3 Experimental Settings
To maximize the capability of these networks, we use the official codes and training configurations provided by different methods to train the models [^3] We have also tried to train all networks with a unified configuration, but find it unreasonable. The performance of networks may vary greatly with different training configurations and optimization strategies, making it difficult to determine a fair unified setting. . Note that all models are trained without using any pre-training strategy (e.g., ×$\times 2$ pre-training for SR) or special tricks (e.g., EMA in SwinIR and TLC in NAFNet) for fair comparison. In addition, we find that different methods may not use exactly the same test sets and the same metrics calculation in their papers to report performance. Therefore, we retest all models based on exactly the same data and calculate PSNR/SSIM using the popular open-source toolbox BasicSR [43].
Table 2: Quantitative results (PSNR/SSIM) of the benchmark. The best and second-best performance results are in bold and underline.
@@
MPRNet 28.90/0.7889 26.88/0.8081 28.48/0.8087 29.71/0.8847 32.66/0.9392 30.96/0.9188 30.29/0.8983 30.43/0.8905 40.34/0.9936
SwinIR 29.07/0.7944 27.47/0.8258 28.56/0.8118 29.88/0.8861 31.66/0.9209 29.41/0.8957 30.05/0.8996 30.45/0.8952 29.14/0.9676
Uformer 27.14/0.7398 25.60/0.7651 28.55/0.8130 29.98/0.8900 33.05/0.9417 30.89/0.9203 27.93/0.8910 24.06/0.8454 33.58/0.9856
Restormer 29.06/0.7934 27.32/0.8199 28.60/0.8130 30.02/0.8898 32.92/0.9398 31.22/0.9226 32.03/0.9237 31.48/0.9054 41.87/0.9938
NAFNet 29.03/0.7919 27.00/0.8112 28.52/0.8098 29.65/0.8840 33.08/0.9419 31.22/0.9237 30.33/0.9097 32.83/0.9135 38.97/0.9928
Figure 4: Visual comparison of Restormer and SwinIR. SwinIR reconstructs significantly more details than Restormer.
@@ 3.4 Benchmark Results
We present the quantitative results for each task on benchmark test sets in Table 2. (Due to space constraints, complete results are provided in Supp.) Several important observations can be made from the benchmark results: 1) Different networks exhibit varying performance on different tasks. For instance, SwinIR performs best on SR but worst on deblurring and dehazing. Uformer excels on denoising and deblurring but performs poorly on deraining and SR. 2) Networks with U-shape and multi-stage architectures demonstrate clear advantages on deblurring and dehazing. 3) MPRNet and NAFNet, which are mainly based on convolution operators, exhibit moderate performance across all tasks without outstanding results. 4) SwinIR, which employs plain architecture and spatial self-attention operators, outperforms other networks by a significant margin on SR. 5) The overall performance of Restormer is outstanding. Except for consistently being weaker than SwinIR on SR, it obtains considerable performance on almost all other tasks.
@@ 3.5 Analysis
We explain the above observations by analyzing the characteristics of different tasks and backbone networks.
@@
The degradation of SR lies in the compression of local information, resulting in a large loss of high-frequency details. Therefore, SR networks often require strong spatial information interaction capability, or even generative capability. The U-shape architecture, which incorporates multiple downsampling operations, intuitively escalates the difficulty of detail reconstruction. In contrast, the plain architecture that maintains feature sizes benefits SR. Besides, window self-attention has demonstrated a superior local fitting ability than convolution [8]. As a result, SwinIR, which is based on a plain structure and employs spatial self-attention operators, exhibits a distinct advantage on SR.
@@
Denoising entails smoothing the image to eliminate high-frequency noise and integrating low-frequency information to reconstruct a clear image. It is conjectured that the high performance of Restormer on denoising can be attributed to its ability to better smooth noise through channel-based processing, akin to operating in the frequency domain. Conversely, SwinIR performs well due to the robust reconstruction ability of its local self-attention.
@@
Deblurring (specifically for motion blur here) involves addressing global motion shifts in the image. As a result, the use of global or multi-scale information is necessary for the deblurring network. Thus, the networks based on the U-shape architecture all perform well on this task. In contrast, SwinIR, which focuses more on local information processing, performs much worse than other networks.
@@
Similar phenomena can be observed for dehazing. Due to the involvement of the depth information in the haze model, the ability to use large-range or even global information is crucial. Besides, dehazing networks are required to handle low-frequency transformations, including alterations in color and contrast, both of which constitute global mappings. Therefore, SwinIR and Uformer, which rely on local spatial information interaction, perform poorly. In contrast, Restormer exhibits exceptional performance.
@@
Deraining is relatively unique in that the rain is unevenly distributed in images, with significant differences between different raindrops and streaks. Thus, there is no clear pattern in the performance of different networks on deraining.
@@
Based on the above results and analysis, we can infer that the acceptable performance of a backbone network on a specific task is predicated on meeting the functional requirements of that task. It is notable that Restormer obtains exceptional task generality. This can be attributed to several factors: 1) The U-shape architecture enables the network to accommodate large-size input. 2) The transposed self-attention allows direct interaction of global information. 3) The presence of depth-wise convolution enables the network to process spatial information effectively. In summary, due to Restormer’s comprehensive functionality, it is capable of meeting the requirements of various tasks.

### 14 Mar 2024 v3

# Chinese Academy of Sciences ${}^{3}$ Lab, Tencent PCG ${}^{4}$hanghai AI Laboratory https://github.com/Andrew0613/X-Restormer

@@ 1 Introduction
Image restoration aims to generate high-quality images from degraded images. In recent years, deep learning has achieved great success in this field, with numerous networks being proposed to address various image restoration tasks. Initially, networks are primarily designed to solve specific restoration tasks and are typically validated only on selected tasks. As deep learning techniques have continued to evolve, there has been an increasing focus on the development of general-purpose networks that can be applied to a broad range of tasks. This trend is particularly evident in the high-level vision field, where new backbone networks are being designed to support multiple tasks [17, 30], including classification, detection and segmentation. For image restoration, although more and more restoration backbones can handle multiple restoration tasks, their task generality is still limited, as illustrated in Footnote 2. For instance, SwinIR [25] achieves state-of-the-art performance on image super-resolution (SR) but falls short on image deblurring and dehazing. Conversely, Restormer [48] performs exceptionally well on image dehazing and deraining but is less effective on image SR. This discrepancy can be attributed to the fact that the characteristics of image degradation vary across different image restoration tasks. While all image restoration tasks involve mapping degraded images to clean images, the requirements for the capability of backbone networks differ depending on specific tasks.
@@
Figure 1: Relative performance difference of different backbone networks on five image restoration tasks [^2] We set the minimum average performance of the networks on test sets in Tab. 2 for the task (i) as the lower bound $P^{(i)}_{lower}$, and set the average performance of X-Restormer for each task as the upper bound $P^{(i)}_{upper}$. The ordinate of each point in the figure with performance $P^{(i)}$ is calculated by $(P^{(i)}-P^{(i)}_{lower})/P^{(i)}_{upper}$.. The existing representative networks exhibit diverse performance on these tasks, while our method presents superior task generality.
@@
Designing a general image restoration backbone network presents a significant challenge. However, the development of such a network holds considerable value, as it has the potential to greatly reduce costs associated with research and application. To achieve this goal, we first conduct a comparative study of mainstream backbone networks on the representative tasks, including image SR, denoising, deblurring, deraining and dehazing. These five tasks are chosen due to the distinct characteristics of their degradation. Five representative backbone networks are selected in the study, including MPRNet [49], Uformer [43], SwinIR [25], Restormer [48] and NAFNet [5]. These five networks encompass classic architectures such as U-shape architecture, plain residual-in-residual architecture and multi-stage progressive architecture. They also employ several common operators, including convolution, spatial self-attention and transposed self-attention [48]. We benchmark the five representative methods on the selected five tasks. The experimental results clearly reflect the performance disparity of different backbone networks on different tasks. We then conduct a detailed analysis of the characteristics of these tasks and these backbone networks to explain the reasons behind the performance differences. Based on the comparative study, we propose that a general backbone network must be highly comprehensive in terms of functionality that meets the diverse needs of various tasks.
@@
It is noteworthy that Restormer stands out in the comparative study, ranking within the top two across all five tasks. This superior performance can be attributed to several key designs. First, Restormer’s U-shape architecture allows it to process large-size inputs, which is crucial for the tasks that deal with large areas of degradation. Then, the network employs transposed self-attention that utilizes channel-wise features as tokens, achieving the information interaction among channels and enabling the mapping with a global receptive field. Additionally, the incorporation of numerous depth-wise convolutions activates the considerable spatial information interaction ability of the network. From a functional perspective, Restormer integrates the key capabilities of the other compared networks, thereby exhibiting commendable task generality in the comparative study. However, the spatial mapping ability of Restormer still appears to be somewhat deficient, as indicated by its quantitatively and qualitatively subpar performance in comparison to SwinIR for SR [^3] In general, models’ SR performance is highly related to the spatial mapping ability..
@@
This inferiority is hypothesized to originate from the inherent challenge of detail reconstruction posed by the U-shape architecture, coupled with the relatively weak spatial mapping capability of depth-wise convolution, particularly when compared to spatial self-attention (i.e., window-based self-attention in SwinIR). To address this limitation, a plausible solution is the introduction of spatial self-attention to Restormer. To achieve this design, we alternately replace half of transposed self-attention blocks with overlapping cross-attention blocks [8], which are proven to have strong spatial information interaction capability, to construct a new network, X-Restormer. Extensive experiments show that this simple modification can significantly enhance the performance of Restormer without increasing the number of parameters. Moreover, our X-Restormer obtains state-of-the-art performance on all five tasks, exhibiting the best task generality.
@@
    We conduct a comparative study by constructing an image restoration benchmark, highlighting the challenges faced by existing image restoration backbone networks in task generality.
@@
    Based on the benchmark results, we perform a detailed analysis of the characteristics of different degradations and networks. We emphasize that the general image restoration backbone network design must meet the functional requirements of diverse tasks.
@@
    By further enhancing the spatial mapping ability of Restormer, we design a preliminary general backbone network, X-Restormer. Without additional parameters, X-Restormer achieves significant performance improvement over existing networks and exhibits superior task generality.
@@ 2 Related Work
Image restoration networks. In the past years, numerous deep networks have been proposed for various image restoration tasks such as image SR [26, 59, 8], denoising [53, 54, 43], deblurring [1, 6], deraining [45, 4, 29] and dehazing [46, 41, 40]. Initially, most deep networks are designed for specific tasks. For instance, Zhang et al. [57] uses the channel attention to construct a CNN-based SR network. Zhang et al. [55] builds a denoising network by inputting multiple downsampled sub-images and a noise level map. Fu et al. [14] employs residual connections to design a deraining network. Cai et al. [3] estimates the transmission rate and inverts the atmospheric scattering model using a CNN to achieve image dehazing. Nah et al. [34] designs a multi-scale network for image deblurring.
Figure 2: Selected five representative image restoration tasks with various degradation.
Recently, more and more methods have been developed to tackle multiple image restoration tasks. Zamir et al. [49] builds a multi-stage CNN for deraining, deblurring and denoising. Wang et al. [43] designs a U-shape Transformer for deraining, deblurring and denoising. Liang et al. [25] implements a Swin Transformer-based network that achieves state-of-the-art performance on SR, denoising and compression artifact reduction. Zamir et al. [48] proposes a novel transposed self-attention to build a U-shape network for deraining, deblurring and denoising. Chen et al. [5] constructs a U-shape CNN for denoising and deblurring. While existing methods have demonstrated some ability to generalize across several restoration tasks, their task generality remains limited.
Difference from the previous network design research. While previous works have proposed networks that excel in various image restoration tasks, their primary focus is on constructing stronger networks to achieve performance breakthroughs on specific tasks. In contrast, this work pays more attention to the task generality of the backbone network, possessing a vision different from previous works. More specifically, our objective is to explore the design principles and directions of general image restoration networks. We are not seeking to create powerful networks for peak performance on a single or some specific tasks, but rather to ensure satisfactory performance across a diverse range of tasks. Regarding the concrete implementation, we do not intend to construct complex network architectures or modules. Our preference, rather, is to enhance task generality through the use of the simplest methodology available.
@@ 3.1 Overview of Image Restoration Tasks
We select five representative tasks for the benchmark experiments. These tasks, exemplified in Fig. 2, are chosen based on two primary reasons. First, they are very common image restoration tasks with widely accepted evaluation schemes. Second, the degradation characteristics of these tasks are diverse and differ greatly from each other. As such, they can provide a robust way to evaluate the task generality of image restoration backbone networks.
@@
where $f_{motion}^{t}(\cdot)$ represents the motion function under different continuous exposure times. This degradation is related to the global motion offset of the image. Therefore, the ability to utilize large-range information and even global information is important for deblurring networks.
@@
where $R$ denotes the additive rain streak, simulated by the physics models, such as [24, 28]. The difference between this degradation and Gaussian noise is that the added $R$ is not evenly distributed on the image and has a correlation with the image content. Complicated rain streaks also places high demands on the complexity of deraining networks.
@@
where $t(\cdot)$ represents the transmission function and $t(I_{GT})$ is associated with the distance from the scene point to the camera. This degradation is intrinsically linked to the depth information within the image. Consequently, the incorporation of global information is important for dehazing networks.
@@ 3.2 Characteristics of Typical Backbone Networks
The architectures of mainstream image restoration networks can be broadly classified into three categories: U-shape encoder-decoder, plain residual-in-residual and multi-stage progressive. Schematic diagrams of these architectures are provided in Supp. The U-shape encoder-decoder architecture performs down-sampling and up-sampling operations on features, enabling networks to handle features of varying scales. This architecture allows networks to accept large-size input, and the effective receptive field of the network expands rapidly with down-sampling. Typical U-shape networks include Uformer [43], Restormer [48]. The multi-stage architecture divides the entire network into several sub-networks and progressively processes features, which are primarily used for image deraining and deblurring. Common networks based on this architecture include MPRNet [49] and HINet [6]. The plain residual-in-residual architecture is composed of several residual groups, each of which consists of several residual blocks. This architecture maintains the original size when processing features, which is favorable for the reconstruction of high-frequency information, but it comes at a high computational cost. Typical networks include RCAN [57] and SwinIR [25].
The core operators for constructing an image restoration network can be mainly categorized into three types: convolution, spatial self-attention and transposed self-attention. These operators are shown in Fig. 3. The convolution calculates a fixed-size filter and processes the entire feature map through a sliding window, which is the major component of many networks, such as RDN [60]. Spatial self-attention is typically implemented as window self-attention in image restoration tasks. It calculates the attention matrix within a fixed window size, generating content-aware weights that are functionally similar to a large kernel dynamic filter. This operator has strong local fitting ability and shows superior advantages on SR and denoising [7]. Transposed self-attention treats the entire feature of each channel as a token to calculate the attention matrix on the channel dimension. This operator directly deals with global features, and when combined with depth-wise convolution, it shows remarkable performance in multiple restoration tasks [48]. The selected five representative backbone networks for the benchmark experiment encompass the abovementioned architectures and core operators, as presented in Tab. 1.
Table 1: Architectures and core operators of the five selected backbone networks.
@@ 3.3 Experimental Settings
To maximize the capability of these networks, we use the official codes and training configurations provided by different methods to train the models [^4] We tried to train all networks with a unified configuration, but find it unreasonable. The performance of networks may vary greatly with different training configurations and optimization strategies, making it difficult to determine a fair unified setting. . Note that all models are trained without using any pre-training strategy (e.g., ×$\times 2$ pre-training for SR) or special tricks (e.g., EMA in SwinIR and TLC in NAFNet) for fair comparison. In addition, we find that different methods may not use exactly the same test sets and the same metrics calculation in their papers to report performance. Therefore, we retest all models based on exactly the same data and calculate metrics using the popular open-source toolbox BasicSR [42].
Table 2: Quantitative results on PSNR(dB) of the benchmark experiments. The best and second-best performance results are in bold and underline.
@@
MPRNet 28.90 26.88 28.48 29.71 32.66 30.96 30.29 30.43 40.34
SwinIR 29.07 27.47 28.56 29.88 31.66 29.41 30.05 30.45 29.14
Uformer 27.14 25.60 28.55 29.98 33.05 30.89 27.93 24.06 33.58
Restormer 29.06 27.32 28.60 30.02 32.92 31.22 32.03 31.48 41.87
NAFNet 29.03 27.00 28.52 29.65 33.08 31.22 30.33 32.83 38.97
@@ 3.4 Benchmark Results
We present the quantitative results of the benchmark experiments in Tab. 2. (Due to space constraints, complete results are provided in Supp.) Several important observations can be made from the results: 1) Different networks exhibit varying performance on different tasks. For instance, SwinIR performs best on SR but worst on deblurring and dehazing. Uformer excels on denoising and deblurring but performs poorly on deraining and SR. 2) Networks with U-shape and multi-stage architectures present clear advantages on deblurring and dehazing. 3) MPRNet and NAFNet, which are mainly based on convolution operators, exhibit moderate performance across all tasks without outstanding results. 4) SwinIR, which employs plain architecture and spatial self-attention operators, outperforms other networks by a significant margin on SR. 5) The overall performance of Restormer is outstanding. Except for consistently being weaker than SwinIR on SR, it obtains considerable performance on almost all other tasks.
@@ 3.5 Analysis
In this section, we explain the above observations by analyzing the characteristics of different tasks and backbone networks.
@@
The degradation of SR lies in the compression of local information, resulting in a large loss of high-frequency details. Therefore, SR networks often require strong spatial information interaction capability, or even generative capability. The U-shape architecture, which incorporates multiple downsampling operations, may undermine the reconstruction of high-frequency information and intuitively escalates the difficulty of detail reconstruction. In contrast, the plain architecture that maintains feature sizes benefits SR. Besides, window self-attention has demonstrated a superior local fitting ability than convolution [8]. As a result, SwinIR, which is based on a plain structure and employs spatial self-attention operators, exhibits a distinct advantage on SR.
@@
Denoising entails smoothing the image to eliminate high-frequency noise and integrating low-frequency information to reconstruct a clear image. This task places no explicit unique requirement for the network, while its performance intuitively benefits from effective spatial information interaction. It is conjectured that the high performance of Restormer on denoising can be attributed to its ability to better smooth noise through channel-wise processing, akin to operating in the frequency domain. In contrast, SwinIR and Uformer perform well due to their robust spatial information interaction ability of the spatial self-attention.
@@
Deblurring (specifically for motion blur here) involves addressing global motion shifts in the image. As a result, the ability to handle large-size inputs and the use of global or multi-scale information are necessary for deblurring networks. Thus, the networks based on the U-shape architecture all perform well on this task. Conversely, SwinIR, which employs the plain architecture and focuses more on local information processing, performs much worse than other networks.
@@
Similar phenomena can be observed for dehazing. Due to the involvement of the depth information in the haze model, the ability to use large-range or even global information is crucial. Besides, dehazing networks are required to handle low-frequency transformations, including alterations in color and contrast, both of which constitute global mappings. Therefore, SwinIR and Uformer, which rely more on local spatial information interaction, perform poorly on this task. On the contrary, Restormer exhibits exceptional performance.
@@
Deraining is relatively unique in that the rain is unevenly distributed in images, with significant differences between different raindrops and streaks. Thus, there is no clear pattern in the performance of different networks on deraining. Nevertheless, networks with higher complexity present better performance.
@@
Based on the above results and analysis, we can infer that the acceptable performance of a backbone network on a specific task is predicated on meeting the functional requirements of that task. It is notable that Restormer obtains exceptional task generality. This can be attributed to several factors: 1) The U-shape architecture enables the network to accommodate large-size input. 2) The transposed self-attention allows direct interaction of global information. 3) The presence of depth-wise convolution enables the network to process spatial information effectively. In summary, due to Restormer’s comprehensive functionality, it is capable of meeting the diverse requirements of different tasks.
Figure 4: Visual and LAM [16] comparisons between Restormer and SwinIR. The LAM results and DI values indicate that Restormer exploits significantly more information than SwinIR. However, SwinIR reconstructs much more details than Restormer.